For a weakly damped harmonic oscillator, the
instantaneous displacement is given by
x(t) = ao exp (-bt) cos (wd t + phy).
Calculate its average energy.
1
Expert's answer
2019-05-15T10:08:24-0400
"E=K+U=0.5m(\frac{dx}{dt})^2+0.5kx^2"

"\frac{dx}{dt}=-a_o \exp{(-bt)} [b \cos{(w_d t + \phi)}+w_d \sin{(w_d t + \phi)}]"

Thus

"K=0.5ma_0^2e^{-2bt} [b^2 \cos^2{(w_d t + \phi)}+w_d^2 \sin^2{(w_d t + \phi)}+bw_d \sin{2(w_d t + \phi)}]"

We have:


"k=mw_d^2"


"U=0.5ma_0^2w_d^2 e^{-2bt} \cos^2{(w_d t + \phi)}""\langle E \rangle=0.5ma_0^2 e^{-2bt} \langle A(t) \rangle"

"A(t)=(b^2+w_d^2) \cos^2{(w_d t + \phi)}+w_d^2 \sin^2{(w_d t + \phi)}+bw_d \sin{2(w_d t + \phi)}"

"\langle A(t) \rangle=0.5((b^2+w_d^2)+w_d^2)"

As for weakly damped harmonic oscillator


"b\ll w_d"

"\langle A(t) \rangle=w_d^2"

"\langle E \rangle=0.5ma_0^2w_d^2 e^{-2bt} =E_0e^{-2bt}"